# Operator Learning with Domain Decomposition for Geometry Generalization in PDE Solving

### v1

@@ Abstract
Neural operators have become increasingly popular in solving partial differential equations (PDEs) due to their superior capability to capture intricate mappings between function spaces over complex domains. However, the data-hungry nature of operator learning inevitably poses a bottleneck for their widespread applications. At the core of the challenge lies the absence of transferability of neural operators to new geometries. To tackle this issue, we propose operator learning with domain decomposition, a local-to-global framework to solve PDEs on arbitrary geometries. Under this framework, we devise an iterative scheme Schwarz Neural Inference (SNI). This scheme allows for partitioning of the problem domain into smaller subdomains, on which local problems can be solved with neural operators, and stitching local solutions to construct a global solution. Additionally, we provide a theoretical analysis of the convergence rate and error bound. We conduct extensive experiments on several representative PDEs with diverse boundary conditions and achieve remarkable geometry generalization compared to alternative methods. These analysis and experiments demonstrate the proposed framework’s potential in addressing challenges related to geometry generalization and data efficiency.
@@ 1 Introduction
Partial differential equation (PDE) solving is of paramount importance in comprehending natural phenomena, optimizing engineering systems, and enabling multidisciplinary applications Evans, (2022). The computational cost associated with traditional PDE solvers Liu and Quek, (2013); Lu et al., (2019) has prompted the exploration of learning-based methods as potential alternatives to overcome these limitations. Neural operators Li et al., 2020b ; Li et al., (2023, 2024); Liu et al., (2023); Hao et al., (2023), as an extension of traditional neural networks, aim to learn mappings between the functional dependencies of PDEs and their corresponding solution spaces. They offer highly accurate approximations to classical numerical PDE solvers while significantly improving computational efficiency. Despite its success, operator learning, as a data-driven approach, encounters the inherent ‘chicken-and-egg’ problem, revealing an interdependence between operator learning and the availability of data. This dilemma arises from the challenge of simultaneously addressing the inefficiency of classical solvers and acquiring an ample amount of data for neural operator training.
@@
Existing works in alleviating the above challenge explore methods for data augmentation. For instance, Lie point symmetry data augmentation (LPSDA) Brandstetter et al., (2022) generates potentially infinitely many new solutions from an existing solution by exploiting symmetries of the differential operator. However, data augmentation techniques only alleviate the data efficiency issue by increasing the diversity and size of the dataset, with the nature of generalization in PDE solving unscrutinized. In this context, we introduce the ‘ladder of generalizability’, which outlines a hierarchical structure comprising three increasingly challenging problem types: (1) coefficient, involving variations in coefficients, source term and boundary conditions while keeping the PDE and domain geometry fixed; (2) geometry, involving changes in the domain geometry while keeping the PDE fixed; and (3) PDE, encompassing the ability to handle various types of PDEs with diverse dependencies. While existing neural operators have shown capabilities in handling diverse geometries through approaches such as geometry parametrization Li et al., (2023) or coordinate representation Hao et al., (2023), they lack the ability to generalize to entirely novel geometries that differ significantly from those present in the training data distribution. The inability to quickly adapt neural operators to unseen geometries without further generating new data hinders the applicability of neural operator learning to real-world problems in industry.
@@
To tackle this challenge, a natural idea is to break down a domain into some basic shapes where neural operator can generalize well. Domain decomposition methods (DDMs) Toselli and Widlund, (2004); Mathew, (2008) provide the suitable tool for this purpose. Thus, we introduce a local-to-global framework to solve PDEs on arbitrary geometries by combining operator learning with DDMs. This framework consists of three parts: (1) Training data generation: creation of random basic shapes and imposition of appropriate boundary conditions on these shapes. This generated data serves as the training set for the neural operator in our framework. (2) Local operator learning: neural operator training to learn solutions on basic shapes. Data augmentation is utilized to enable the neural operator to capture the intricate details and variations within these shapes. (3) Schwarz neural inference (SNI): a three-step algorithm for inference. Firstly, the computational domain is partitioned into smaller subdomains. Then, the learned operator is applied within each subdomain to obtain the local solution. Finally, an iterative process of stitching and updating the global solution is performed using additive Schwarz methods.
By combining these three components, our local-to-global framework enables the accurate PDE solving on arbitrary geometries. The training data generation ensures the availability of diverse and representative data, the local operator learning captures local behavior effectively, and the SNI algorithm handles the partitioning and integration of solutions across the entire computational domain.
@@
  • We introduce a local-to-global framework that combines operator learning with DDMs. This work represents the first attempt, to our best knowledge, in directly tackling the inherent ‘chicken-and-egg’ problem through the lens of geometry generalization for operator learning.
  • We propose a novel data generation scheme and an inference algorithm SNI.
  • We theoretically analyze the convergence and the error bound of the proposed algorithm. Through comprehensive experiments, we empirically validate the effectiveness of our framework on generalizing to new geometries.
@@ 2.1 Problem formulation
where $\mathcal{L}$ is a self-adjoint and coercive elliptic partial differential operator Mathew, (2008) and $\Gamma_{D}\cup\Gamma_{N}=\partial\Omega$ denotes Dirichlet and Neumann boundary, respectively. We assume all the domains $\Omega$ are bounded orientable manifolds embedded in some ambient Euclidean space $\mathbb{R}^{n}$ Li et al., (2023). Later we will extend our method to handle time-dependent parabolic equations.
@@ 2.2 Domain decomposition methods
Domain decomposition methods (DDMs) solve Eq. (1) by decomposing domain into subdomains and iteratively solve a coupled system of equations on each subdomain. An overlapping decomposition of $\Omega$ is a collection of open subregions $\{\Omega_{k}\}_{k=1}^{K}$, $\Omega_{k}\subseteq\Omega$ for $k=1,\dots,K$ such that $\bigcup_{k=1}^{K}\Omega_{k}=\Omega$. We denote $V$ and $\{V_{k}\}_{k=1}^{K}$ to be finite element space associated with domain $\Omega$ and $\{\Omega_{k}\}_{k=1}^{K}$. We can define restriction operators $\{R_{k}:V\rightarrow V_{k}\}_{k=1}^{K}$ restricting functions on $\Omega$ to $\{\Omega_{k}\}_{k=1}^{K}$ and extension operators $\{R^{\intercal}_{k}:V_{k}\rightarrow V\}_{k=1}^{K}$ extending functions on $\{\Omega_{k}\}_{k=1}^{K}$ to $\Omega$ by zero.
In the subsequent discussion, we revisit the idea of additive Schwarz method (ASM) in DDMs for overlapping decomposition. The additive Schwarz-Richardson iteration Mathew, (2008) has the following form:
@@
We denote the local operator $\mathcal{S}_{k}:(u^{n},u_{D},g)\mapsto w_{k}^{n+1}$. Note that the first two boundary conditions in Eq. (3) is the boundary condition on the global boundary part of $\partial\Omega_{k}$ and is not updated during iteration. The last boundary condition is along the artificial boundary created by decomposition and the value is updated through iteration. Hence $\{\mathcal{S}_{k}\}_{k=1}^{K}$ can be considered as a single-input operator when the global boundary condition and decomposition are determined. This iterative process can be shown to converge for FEMs under mild assumption on properties of equation and decomposition. Please refer to Appendix A.1 for more details.
@@ 3 Operator Learning with Domain Decomposition
In order to solve PDE on arbitrary geometry with neural operator, a natural idea is to decompose domain into a prescribed family of building blocks (basic shapes) since it is not feasible to explicitly consider arbitrary shapes during training process. The idea of marrying operator learning with domain decomposition is to train a neural operator to solve local problems on these shapes and stitch local solutions together to get a global solution, corresponding to the training and inference stages, respectively. An illustration of the proposed framework is presented in Figure 1. A detailed implementation will be discussed in the following subsections.
@@ 3.1 Training Data Generation
Choice of basic shapes. The selection of basic shapes cannot be arbitrary due to the requirement of ensuring the neural operator’s capability in solving local problems across a wide range of shapes. To address this issue, we need to specify a probability space $(\mathcal{P},\mu)$ where $\mu$ denote the probability distribution over $\mathcal{P}$. Moreover, two necessary criteria should be set forth for basic shape generation: (1) sampling feasibility: it should be tractable to sample from $\mu$ and solve boundary value problems on shapes in $\mathcal{P}$ in order to generate data for operator learning; (2) complete coverage: basic shapes in $\mathcal{P}$ should be flexible to cover any shape of domain.
For implementation, we focus on $\Omega\subseteq\mathbb{R}^{2}$. We propose to use the space $\mathcal{P}_{s}(n)$ of simple polygons with at most $n$ vertices (i.e. planar polygon without self-intersection and holes) uniformly bounded by a compact region in $\mathbb{R}^{2}$. Simple polygons are Lipschitz domains with straightforward sampling method Auery and Heldz, (2019) and flexible enough to constitute any discretized planar domain Preparata and Shamos, (2012). We note, however, that this is not the only choice of these basic shapes. We could equally use convex polygons, star-shaped polygons, etc. as long as the two aforementioned criteria are satisfied.
Imposing boundary conditions. The imposition of boundary conditions presents two complications: (1) Types of boundary conditions. Neumann boundary conditions in Eq. (1) will inevitably result in mixed boundary conditions in local subdomains. To generate solutions with mixed boundary conditions, we randomly divide the boundary of a basic shape into two connected components, representing the Dirichlet and Neumann boundaries, respectively. During inference, we have to carefully set hyperparameters for decomposition to make sure boundary of subdomains have at most two connected components for Dirichlet and Neumann boundaries. (2) Functional range of boundary conditions. In general, the inference process for subdomains will encounter arbitrary ranges in boundary conditions. However, it is practically infeasible to train the neural operator to handle unbounded boundary values. Instead, we generate random functions with values normalized within a bounded range for both boundary conditions and other input functions such as coefficient fields and source terms. We will handle this complication with symmetries of PDEs during inference.
@@ 3.2 Local Operator Learning
Choice of neural operator architecture. Our framework is orthogonal to the choice of neural operator architecture as long as the architecture can accommodate flexible input/output formats and possesses sufficient expressive power to solve local problem with randomly varying domain and input functions. For implementation, we adopt GNOT Hao et al., (2023) which is a highly flexible transformer-based neural operator. We note that, however, training neural operator on highly varying geometries presents challenges to both design of architectures and training schemes.
Data augmentation. To enhance the generalization capabilities of the neural operator, Lie point symmetry data augmentation (LPSDA) Brandstetter et al., (2022) can be naturally applied to local solutions during training. Examples of such transformations are rotation and scaling. It is crucial to appropriately extend these transformations to boundary conditions and other input functions, taking into account the symmetries inherent in the PDEs. Please refer to Appendix A.5 for a detailed discussion.
3.3 Schwarz Neural Inference
@@
0:  Domain $\Omega$; Global Boundary Condition $B$; Other input functions $H$; Number of Subdomains $K$; Depth of Extension $d$; Local Operator $\mathcal{G}^{\dagger}$; Step Size $\tau$; Convergence Criterion $C$;
@@
4:     update local boundary condition $\{B_{k}^{n}\}_{k=1}^{K}$ by global boundary condition $B$ and last-step global solution $u^{n}$;
5:     obtain the preprocessing $\{T_{k}\}_{k=1}^{K}$ and postprocessing transformations $\{\tilde{T}_{k}\}_{k=1}^{K}$;
6:     inference on each subdomain using local operator: $\tilde{w}^{n+1}_{k}=\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}(\Omega_{k},B_{k}^{n})$;
7:     extend local solution: $w^{n+1}=\sum_{k=1}^{K}R_{k}^{\intercal}\tilde{w}^{n+1}_{k}+(I-R_{k}^{\intercal}R_{k})u^{n}$;
8:     update global solution: $u^{n+1}=(1-\tau K)u^{n}+\tau w^{n+1}$;
9:     $n=n+1$;
10:  end while
11:  return  $u^{n}$;
Decomposition into overlapping subdomains. In general, there is no natural methods to decompose an arbitrary domain into desired shapes, and here we adopt the common practice in DDM literature Mathew, (2008). We assume there exists a pre-defined triangulation $\mathcal{T}_{h}(\Omega)$ of the domain $\Omega$, and a graph can be constructed to represent the connectivity of this triangulation. A graph partition algorithm such as METIS Karypis and Kumar, (1997) is then employed to partition this graph into $K$ non-overlapping connected subgraphs with index sets ~$\tilde{\mathcal{I}}_{1},\dots,\tilde{\mathcal{I}}_{K}$. To achieve an overlapping decomposition, each subgraph is then extended iteratively by including neighboring vertices for $d$ iterations. This process generates index sets $\mathcal{I}_{1},\dots,\mathcal{I}_{K}$ that, together with the original mesh, form an overlapping decomposition denoted as $\{\Omega_{k}\}_{k=1}^{K}$. An intuitive illustration of this process is depicted in Figure 1.

### v2

# arXiv:2504.00510v2[cs.LG] 27 Feb 2026

@@ Abstract
Neural operators have become increasingly popular in solving partial differential equations (PDEs) due to their superior capability to capture intricate mappings between function spaces over complex domains. However, the data-hungry nature of operator learning inevitably poses a bottleneck for their widespread applications. At the core of the challenge lies the absence of transferability of neural operators to new geometries. To tackle this issue, we propose operator learning with domain decomposition, a local-to-global framework to solve PDEs on arbitrary geometries. Under this framework, we devise an iterative scheme Schwarz Neural Inference (SNI). This scheme allows for partitioning of the problem domain into smaller subdomains, on which local problems can be solved with neural operators, and stitching local solutions to construct a global solution. Additionally, we provide a theoretical analysis of the convergence rate and error bound. We conduct extensive experiments on several representative linear and nonlinear PDEs with diverse boundary conditions and achieve remarkable geometry generalization compared to alternative methods. These analysis and experiments demonstrate the proposed framework’s potential in addressing challenges related to geometry generalization and data efficiency. The code is publicly available at this repository: https://github.com/questionstorer/sni.
@@ 1 Introduction
Partial differential equation (PDE) solving is of paramount importance in comprehending natural phenomena, optimizing engineering systems, and enabling multidisciplinary applications (Evans, 2022). The computational cost associated with traditional PDE solvers (Liu and Quek, 2013; Lu et al., 2019) has prompted the exploration of learning-based methods as potential alternatives to overcome these limitations. Neural operators (Li et al., 2020b; 2023; 2024; Liu et al., 2023; Hao et al., 2023), as an extension of traditional neural networks, aim to learn mappings between the functional dependencies of PDEs and their corresponding solution spaces. They offer highly accurate approximations to classical numerical PDE solvers while significantly improving computational efficiency. Despite its success, operator learning, as a data-driven approach, encounters the inherent ‘chicken-and-egg’ problem, revealing an interdependence between operator learning and the availability of data. This dilemma arises from the challenge of simultaneously addressing the inefficiency of classical solvers and acquiring an ample amount of data for neural operator training.
@@
Existing works in alleviating the above challenges explore symmetries of PDEs. Lie point symmetry data augmentation (LPSDA) (Brandstetter et al., 2022) generates potentially infinitely many new solutions of PDE from existing solution by exploiting symmetries of differential operator defining the PDE. Subsequent work (Mialon et al., 2023) applies LPSDA for self-supervised learning. However, LPSDA only partially alleviate the problem in data efficiency and the problem of how to quickly generalize to new geometry is untouched. While existing neural operators have shown capabilities in handling diverse geometries through approaches such as geometry parametrization (Li et al., 2023) or coordinate representation (Hao et al., 2023), they lack the ability to generalize to entirely novel geometries that differ significantly from those present in the training data distribution. The inability to quickly adapt neural operators to unseen geometries without further generating new data hinders the applicability of neural operator learning to real-world problems in industry.
@@
To tackle this challenge, a natural idea is to break down a domain into some basic shapes where neural operator can generalize well. Domain decomposition methods (DDMs) (Toselli and Widlund, 2004; Mathew, 2008) provide the suitable tool for this purpose. Related efforts such as Mao et al. (2024) have combined operator learning with DDMs on uniform grids to accelerate classical methods. In contrast, our work aims to extend this paradigm to arbitrary geometries through a local-to-global framework. This framework consists of three parts: (1) Training data generation: creation of random basic shapes and imposition of appropriate boundary conditions on these shapes. This generated data serves as the training set for the neural operator in our framework. (2) Local operator learning: neural operator training to learn solutions on basic shapes. Data augmentation based on symmetries of PDEs is utilized to enable the neural operator to capture the intricate details and variations within these shapes. (3) Schwarz neural inference (SNI): a three-step algorithm for inference. Firstly, the computational domain is partitioned into smaller subdomains. Then, the learned operator is applied within each subdomain to obtain the local solution. Finally, an iterative process of stitching and updating the global solution is performed using additive Schwarz methods.
@@
  • We introduce a local-to-global framework that integrates operator learning with domain decomposition methods as an attempt in tackling the geometry generalization challenge in operator learning.
  • We design a novel data generation scheme that leverages random shape generation and symmetries of PDEs to train local neural operators for solving PDEs on basic shapes.
  • We propose an iterative inference algorithm, SNI, built upon a trained local neural operator to obtain solutions on arbitrary geometries. We theoretically analyze the convergence and the error bound of the algorithm for a wide range of elliptic PDEs. Through comprehensive experiments, we empirically validate the effectiveness of our framework on generalizing to new geometries for both linear and nonlinear PDEs.
@@ 2.1 Problem formulation
where $\mathcal{L}$ is a partial differential operator and $\Gamma_{D}\cup\Gamma_{N}=\partial\Omega$ denotes Dirichlet and Neumann boundary, respectively. We assume all the domains $\Omega$ are bounded orientable manifolds embedded in some ambient Euclidean space $\mathbb{R}^{n}$ (Li et al., 2023). Later we will extend our method to handle time-dependent equations.
@@ 2.2 Domain decomposition methods
Domain decomposition methods (DDMs) solve Eq. 1 by decomposing domain into subdomains and iteratively solve a coupled system of equations on each subdomain. An overlapping decomposition of $\Omega$ is a collection of open subregions $\{\Omega_{k}\}_{k=1}^{K}$, $\Omega_{k}\subseteq\Omega$ for $k=1,\dots,K$ such that $\bigcup_{k=1}^{K}\Omega_{k}=\Omega$. We denote $V$ and $\{V_{k}\}_{k=1}^{K}$ to be finite element space associated with domain $\Omega$ and $\{\Omega_{k}\}_{k=1}^{K}$. We can define restriction operators $\{R_{k}:V\rightarrow V_{k}\}_{k=1}^{K}$ restricting functions on $\Omega$ to $\{\Omega_{k}\}_{k=1}^{K}$ and extension operators $\{R^{\intercal}_{k}:V_{k}\rightarrow V\}_{k=1}^{K}$ extending functions on $\{\Omega_{k}\}_{k=1}^{K}$ to $\Omega$ by zero.
In the subsequent discussion, we revisit the idea of additive Schwarz method (ASM) in DDMs for overlapping decomposition. The additive Schwarz-Richardson iteration (Mathew, 2008) has the following form:
@@
We denote the local operator $\mathcal{S}_{k}:(u^{n},u_{D},g)\mapsto w_{k}^{n+1}$. Note that the first two boundary conditions in Eq. 3 is the boundary condition on the global boundary part of $\partial\Omega_{k}$ and is not updated during iteration. The last boundary condition is along the artificial boundary created by decomposition and the value is updated through iteration. Hence $\{\mathcal{S}_{k}\}_{k=1}^{K}$ can be considered as a single-input operator when the global boundary condition and decomposition are determined. This iterative process can be shown to converge for FEMs under mild assumption on properties of equation and decomposition. Please refer to Appendix A for more details.
@@ 3 Operator Learning with Domain Decomposition
In order to solve PDE on arbitrary geometry with neural operator, a natural idea is to decompose domain into a prescribed family of building blocks (basic shapes) since it is not feasible to explicitly consider arbitrary shapes during training stage. For that purpose, we propose to train a neural operator to solve local problems on basic shapes and stitch local solutions together to get a global solution. An illustration of the proposed framework is presented in Figure 1. A detailed implementation will be discussed in the following subsections.
@@ 3.1 Training Data Generation
Choice of basic shapes. The selection of basic shapes cannot be arbitrary due to the requirement of ensuring the neural operator’s capability in solving local problems across a wide range of shapes. Moreover, three necessary criteria should be set forth for basic shape generation: (1) sampling feasibility: it should be tractable to sample from and solve boundary value problems on basic shapes in $\mathcal{P}$. (2) complete coverage: basic shapes in $\mathcal{P}$ should be flexible to cover any shape of domain.
For implementation, we focus on $\Omega\subseteq\mathbb{R}^{2}$. We propose to use the space $\mathcal{P}_{s}(n)$ of simple polygons with at most $n$ vertices (i.e. planar polygon without self-intersection and holes) uniformly bounded by a compact region in $\mathbb{R}^{2}$. Simple polygons are Lipschitz domains with straightforward sampling method (Auery and Heldz, 2019) and flexible enough to constitute any discretized planar domain (Preparata and Shamos, 2012). We note, however, that this is not the only choice of these basic shapes. We could equally use convex polygons, star-shaped polygons, etc. as long as the two aforementioned criteria are satisfied.
Imposing boundary conditions. The imposition of boundary conditions presents two complications: (1) Types of boundary conditions. Neumann boundary conditions in Eq. 1 will inevitably result in mixed boundary conditions in local subdomains. To generate solutions with mixed boundary conditions, we randomly divide the boundary of a basic shape into two connected components, representing the Dirichlet and Neumann boundaries, respectively. During inference, we have to carefully set hyperparameters for decomposition to make sure boundary of subdomains have at most two connected components for Dirichlet and Neumann boundaries. (2) Functional range of boundary conditions. In general, the inference process for subdomains will encounter arbitrary ranges in boundary conditions. However, it is practically infeasible to train the neural operator to handle unbounded boundary values. Instead, we generate random functions with values normalized within a bounded range for both boundary conditions and other input functions such as coefficient fields and source terms. We will handle this complication with symmetries of PDEs during inference.
@@ 3.2 Local Operator Learning
Choice of neural operator architecture. Our framework is orthogonal to the choice of neural operator architecture as long as the architecture can accommodate flexible input/output formats and possesses sufficient expressive power to solve local problem with randomly varying domain and input functions. For implementation, we adopt GNOT (Hao et al., 2023) which is a highly flexible transformer-based neural operator. We note that, however, training neural operator on highly varying geometries presents challenges to both design of architectures and training schemes.
Data augmentation. To enhance the generalization capabilities of the neural operator, Lie point symmetry data augmentation (LPSDA) (Brandstetter et al., 2022) can be naturally applied to local solutions during training. Examples of such transformations are rotation and scaling. It is crucial to appropriately extend these transformations to boundary conditions and other input functions, taking into account the symmetries inherent in the PDEs. Please refer to Appendix E for a detailed discussion.
@@
0: Domain $\Omega$; Global boundary data $B$ (Dirichlet/Neumann); Other input functions $H$ (e.g., coefficient fields, source terms, PDE parameters); Number of subdomains $K$; Depth of extension $d$; Local operator $\mathcal{G}^{\dagger}$; Step size $\tau$; Convergence criterion $C$;
@@
4:  Form local inputs $\{(\Omega_{k},B_{k}^{n},H_{k})\}_{k=1}^{K}$ from $\Omega$, $B$, $H$ and the last-step solution $u^{n}$;
5:  Obtain preprocessing transforms $\{T_{k}\}_{k=1}^{K}$ and postprocessing transforms $\{\tilde{T}_{k}\}_{k=1}^{K}$;
6:  Local inference on each subdomain: $\hat{w}^{n+1}_{k}=\tilde{T}_{k}\big(\mathcal{G}^{\dagger}\big(T_{k}(\Omega_{k},B_{k}^{n},H_{k})\big)\big)$;
7:  Global update (additive Schwarz form): $u^{n+1}=u^{n}+\tau\sum_{k=1}^{K}R_{k}^{\intercal}\big(\hat{w}^{n+1}_{k}-R_{k}u^{n}\big)$;
8:  $n=n+1$;
9:end while
10:return $u^{n}$;
3.3 Schwarz Neural Inference
At each iteration $n$, SNI assembles, for every subdomain $\Omega_{k}$, a local boundary input $B_{k}^{n}$ by combining the prescribed global boundary data on $\partial\Omega_{k}\cap\partial\Omega$ and the interface Dirichlet data from the previous iterate on $\partial\Omega_{k}\cap\Omega$. For PDEs with additional input functions or parameters (e.g., coefficient fields and source terms), we denote by $H_{k}$ their restriction to $\Omega_{k}$. In Algorithm 1, $\hat{w}_{k}^{n+1}$ denotes the post-processed local solution on $\Omega_{k}$.
Decomposition into overlapping subdomains. In general, there is no canonical way to decompose an arbitrary domain into prescribed shapes. Here we adopt a common practice in the DDM literature (Mathew, 2008). We assume there exists a pre-defined triangulation $\mathcal{T}_{h}(\Omega)$ of the domain $\Omega$, and a graph can be constructed to represent the connectivity of this triangulation. A graph partition algorithm such as METIS (Karypis and Kumar, 1997) is then employed to partition this graph into $K$ non-overlapping connected subgraphs with index sets $\tilde{\mathcal{I}}_{1},\dots,\tilde{\mathcal{I}}_{K}$. To achieve an overlapping decomposition, each subgraph is then extended iteratively by including neighboring vertices for $d$ iterations. This process generates index sets $\mathcal{I}_{1},\dots,\mathcal{I}_{K}$ that, together with the original mesh, form an overlapping decomposition denoted as $\{\Omega_{k}\}_{k=1}^{K}$. An intuitive illustration of this process is depicted in Figure 1.